# Environment-Specific People

### v1

@@ 3.3 HPM translation module
In the remainder of the pipeline, we use diffusion models [17]. In particular, we leverage the large pre-trained latent diffusion model that is called Stable Diffusion [41] as our image prior. Diffusion models are well-known for their capacity to generate highly detailed images. They do so gradually through the denoising diffusion process in $T$ timesteps. During training, Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ gets added to the original sample $\mathbf{x}_{0}$, step by step, and a model similar to UNet [18] is trained to predict the noise $\epsilon$ residuals, at timestep $t$. For latent diffusion models where the denoising process is performed in the latent space rather than in pixel space, $\mathbf{z}_{t}$ marks the corresponding latent representation of a noisy sample $\mathbf{x}_{t}$ at timestep $t$. The LDM objective is then formulated as follows:
@@
Given the large success of LDMs, new mechanisms have appeared on how to control their image generation process further. In this paper we use the approach called ControlNet from [52]. The main idea behind ControlNet is to create a trainable copy of a pre-trained LDM, attach it to the LDM and in turn train it without modifying the original model (zero-convolutions). In this way, generated images retain the quality from the original LDM which remains unchanged. The copy is additionally conditioned on images $\mathbf{c}_{\text{f}}$ that are encoded into its latent space by a small learnt neural network and the feature maps of the encoder of the LDM copy are added to the feature maps of the original LDM so as to transfer the original conditioning signal. Including the text conditioning $\mathbf{c}_{t}$, objective fuction for a conditional LDM is now equal to:
@@
We attach a ControlNet module to the Stable Diffusion model and train it on approx. 23K images where the input conditioning are RGB HPM images while the output are the corresponding images of real people in the wild. This model is later used as a translator for our custom low resolution RGB HPM images that are padded and then upscaled from 32x64 to 512x512 during training. The output is upscaled from 200x200 to 512x512 during training.
@@ 3.4 Inpainting and super-resolution
Human body is highly complex due to its high flexibility that stems from the many degrees of freedom that it has. Also, one body may be clothed in many different types of outfits. For this reason, there exists a large semantic gap between the body and its binary segmentation mask. This gap makes it especially difficult to inpaint human bodies into photographs, when only the ambiguous binary mask is available and the body is to be inpainted into the specified region. To this end, at inference time, we condition the inpainting LDM on a semantic representation of the body and clothing (HPM) via our pre-trained HPM translation module (ControlNet). In this way, we are able to transfer the background from the original image into the resulting image using the inpainting model while predicting the HPM-compatible humans at the same time. Additionally, by using a weighting scheme where the impact of the HPM translator is reduced, it is possible to generate images of very high quality where the quality is similar to that of the Stable Diffusion-generated imagery. This is done by reducing the weight $\beta$ (controlNet strength) that is used for the multiplicaton of the ControlNet’s feature maps prior to their addition with the feature maps of the inpainting model. However, $\beta$ should not be reduced drastically as the benefits of the guidance signal will become undone. We find that setting $\beta$ to $0.5$ works well.
Figure 3: Fixing the context (rows F) generates more uniform clothing than varying the context embeddings (rows R). This indicates there is a link between the context and the environment that has been learned by our context-aware HPM generator.
Figure 4: Background Reconstruction: VAE reconstructions are shown in rows 1 (GT) and 2 (VAE); fixing the context vector in the StyleGAN generator gives matching results (row F) with GT/VAE, while varying the context leads to random predictions (row R).
Figure 5: SOTA Comparison: the environment where the humans are to be inpainted is shown in the first row; results with StyleGAN-Human [12] are in row 2; ControlNet [52] results are in row 3. Our method generates more appropriate, environment-specific, people.
Output $\psi=0.5$ $\psi=0.6$ $\psi=0.7$ $\psi=0.8$ $\psi=0.9$ $\psi=1$
HPM 15.73 14.59 13.53 12.64 11.96 11.49
Context 112.45 86.33 67.55 54.77 46.29 41.08
Table 1: Truncation Trick: the table shows FIDnormal-$\downarrow$ scores for varying $\psi$ across 10K randomly selected contexts. The optimal value of $\psi$ is equal to $1.0$ for both the HPM and the Context generator.
Method $\beta=0$ $\beta=0.25$ $\beta=0.5$ $\beta=0.75$ $\beta=1$
Baseline - - 78.7 86.7 90.0
Ours 42.3 22.6 21.0 19.8 20.2
Table 2: ControlNet Strength: the table shows FIDnormal-$\downarrow$ scores for varying $\beta$ across 10K images. Our method outperforms SOTA.
4 Experiments
4.1 Data
Our method utilizes a training dataset comprising context images, images with humans, skeletal pose information, and machine-annotated human body parsing or semantic segmentation maps. For training with the AHP dataset [57], we extract context images by using segmentation masks to remove people from the images. Skeleton images with pose data are generated using OpenPose [7, 6] on these human-included context images. During inference, only a context image and pose information are required to generate a full-body person image with context-aware appearance.
@@ C Evaluation PSNR $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Table 3: Quantitative Analysis: the results are calculated over 10K contexts. C denotes whether the context vector is fixed (F) or random (R). Overall, fixing the context vector leads to the predictions that are closer to the ground truth, especially in background reconstruction. Therefore, the analysis confirms that our method is context-aware (PSNR, MSE, LPIPS) and that it generates meaningful parsing masks and background contexts (FID, KID). The performance of the standalone VAE module used to generate the blurry contexts (rows 2 and 5) is shown in the last row.
The AHP dataset is compiled from large-scale instance segmentation and detection datasets, including COCO [29], VOC [10], SBD [16], and Objects365 [46]. It underwent filtering and post-processing to yield 56,599 images with parsing maps and binary human masks, which are integral to our method. However, the raw AHP dataset’s varied resolution and quality, along with inconsistent human pixel percentages, lead to suboptimal results. Additionally, the provided parsing maps sometimes have disconnections and holes.
@@
To mitigate these issues, we further refined the AHP images. Images were selected based on human size, with a minimum dimension of 180 pixels in height and 90 pixels in width, and standardized to a resolution of $200\times 200$. We also filtered out images with mismatched binary masks and parsing maps to prevent background bleed into human images. To address occlusions in parsing maps, we added an additional segmentation category for accessories and discarded images with disconnected parsing map parts. Our final dataset consists of about 23K images, each paired with its corresponding parsing map and OpenPose keypoints, ensuring higher consistency and quality for training.
Figure 6: Super-Resolution: lowering the impact of the fine-tuned LDM copy through $\beta$, higher quality may be achieved. When $\beta$ is too low, humans are no longer generated (column 4).
4.2 Comparison with state-of-the-art
@@
Qualitative/quantitative analysis: Figure 5 contains our results in comparison with an unconditional GAN-based SOTA method (StyleGAN-Human [12]) as well as a diffusion-based SOTA method that has been trained using similar 2D pose conditioning as ours (ControlNet [52]) and high-resolution images. In the former, humans are generated at 1024x1024, the background is subsequently removed using rembg [2] and humans are then pasted on top of the context image, at random. In the latter, we use Stable Diffusion for inpainting where the rectangular masks are passed to the inpainting model as the ControlNet model does not provide any information about the location of the body. Our method generates more natural-looking, environment-specific humans the clothing semantics and the overall color harmonization are considered. Additionally, our quantitative analysis in Table 2 shows how our method outperforms ControlNet.
In Figure 6, we gradually decrease the impact of our fine-tuned ControlNet model through $\beta$. When $\beta$ equals 00, our system is equivalent to the vanilla Stable Diffusion inpainting model, using a binary mask. It is observed that, in this case, the output does not contain people in the desired regions. This phenomenon occurs because of a large semantic gap that exists between the complex human body and the binary inpainting mask. This experiment is to show how current, diffusion-based, state-of-the-art inpainting models, do not have the capacity to inpaint humans into existing photographs without additional guidance (HPM in our case). Numerical results are included in Table 2.
Figure 7: Pose is fixed to show how our method generates plausible clothing independent of the pose and depending on the context.
Figure 8: Results: ESP outputs a wide range of people in clothing suitable for diverse contexts. More results are included in the appendix.
@@
Perceptual evaluation: To evaluate the appropriateness of our generated humans versus state-of-the-art, we use BLIP-2 [27] to caption a set of context images and group them by the environment type. We then select 20 prominent environment groups and randomly select one sample for each. We compete against the current state-of-the-art, i.e. a pose-conditioned LDM from [52]. 16 people were attending the short study and the results show that our method was preferred 61,875% of the time. 3.4375% of the time, the user was undecided.
@@ 4.3 Evaluation and ablation studies
Context embeddings: Figure 4 shows the blurry background reconstruction results of our (1) VAE, and (2) contextual HPM StyleGAN generator. Both predict images that are similar to the GT. StyleGAN generator’s predictions are of lower quality than the VAE. This is also noticeable in the quantitative analysis. Nevertheless, we find that these reconstructions are meaningful as they also lead to the generation of environment-inspired clothing semantics.
In Figure 3, a difference between the fixed and random context embeddings is shown in the generator. Fixing the context (rows F) generates more uniform clothing. Here, each sample is generated with a different random part of the contextual style vector, while the context embeddings are fixed. On the contrary, when the context embeddings are not fixed, our model generates more varied clothing styles (rows R). This indicates there is a link between the context and the environment, as expected.
Pose conditioning: Previously, it had been shown that pose plays an important role in the synthesis of scene-compatible humans [5]. Here we argue that the clothing semantics may also be inferred from the scene context alone. Therefore, we fix the pose in Fig. 7 to show how our method generates plausible clothing independent of the pose and depending on the image context. More samples where the original pose is used are shown in Fig. 8. For examples where the conditioning is shown, see Fig. 1.
Inpainting: Changing the impact of our ControlNet model $\beta$ may increase image quality (Figure 6). This is due to the low resolution of images that we use to train the model. When $\beta$ is lowered, the original Stable Diffusion model that was trained on high-quality images has more impact on the final result, and therefore images of higher quality are generated. However, if $\beta$ is too low, generated humans might no longer respect the input pose or they might not be generated, as discussed above.
@@
Quantitative analysis: To determine the effectiveness of our HPM generator, we compare it to GT data with a fixed context C (denoted as F) and with a random context (R) in Table 3. The results demonstrate that our method is able to generate meaningful parsing masks and background images when the original data distribution is considered (FID, KID). It is also shown, quantitatively, that the background reconstruction errors (PSNR, MSE, LPIPS) are considerably lower when the context vector is fixed, indicating that our method generates context-aware predictions. Our reconstructed backgrounds (row 2) are similar to the VAE-reconstructed backgrounds (row 7). The VAE module provides a guiding vector that makes it possible for our HPM model to generate backgrounds that look similar to the real backgrounds (row 3). Since the HPM and the background generator share all of the parameters, we conclude that our HPM generation is context-aware. Numerical results for inpainting including the corresponding ablation of $\beta$ can be found in Table 2.
5 Discussion
@@
Our results, as visually compelling as they are, are biased. The biases observed — such as the lack of racial diversity in certain contexts, the uniformity of skin color in scenarios like handshake interactions, the socioeconomic implications of correlating impoverished environments with darker-skinned individuals, and the gender imbalances skewed towards male representations — are a reflection of the training data’s limitations. Therefore, it is imperative to take conscious steps to mitigate these biases which includes diversifying the training data.
@@ 6 Conclusion
In this paper, we have demonstrated how full-body human generation and subsequent inpainting can be made context-aware, i.e., our generated humans are in sync with the environment when their clothing semantics and overall color harmonization are considered. Our main contribution is conditioning the image generation process on scene context to generate environment-specific people in clothing that is suitable for inpainting into a given environment. Our results demonstrate that our method produces synthetic imagery of plausible contextualized people in scenes, outperforming state-of-the-art methods on the task of contextual full-body generation.
Acknowledgement: This project has received funding from the EU Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 860768. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH. While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Appendix
Appendix A Model architecture
Scene representation.
We extract the contextual scene representation vectors from the latent space of a variational auto-encoder (VAE). Here, the encoder consists of $5$ blocks of layers, where each block is comprised of one convolutional, batch normalization, and a leaky ReLU layer. The bottleneck layer is $512$-dimensional, and it contains our context code that is used as a conditioning in the StyleGAN generators for human parsing maps. The decoder mirrors the encoder, and its $5$ blocks are comprised of upconvolution, batch normalization, and a leaky ReLu layer. In the last block, we add an additional convolution and a tanh activation.
The architecture of the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
Appendix B Training details
Scene representation.
We train the scene VAE on the AHP dataset [57] with a batch size of $64$ for approximately $500$ epochs and with a learning rate of $0.0002$. We use Adam optimizer with betas set to $0.5$ and $0.9$. The weights for the reconstruction and KL divergence loss terms are set to $\lambda_{1}=1$ and $\lambda_{2}=10$ (see Equation 1. in the main paper). After the training procedure has been completed, we freeze the VAE’s parameters and use its latent space as the context embedding to condition the contextual human parsing map generators.
The training procedure for the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
Figure 9: Additional results for different contexts.
Appendix C Additional Results
Figure 9 contains more results for various environments.

### v2

@@ 3.3 HPM translation module
In the remainder of the pipeline, we use diffusion models [18]. In particular, we leverage the large pre-trained latent diffusion model Stable Diffusion [43] as our image prior. Diffusion models are well-known for their capacity to generate highly detailed images. They do this gradually, through the denoising diffusion process in $T$ timesteps. During training, Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ gets added to the original sample $\mathbf{x}_{0}$, step by step, and a UNet-like model [19] is trained to predict the noise $\epsilon$ residuals, at timestep $t$. For latent diffusion models (LDM), where the denoising process is performed in the latent space rather than in pixel space, $\mathbf{z}_{t}$ marks the corresponding latent representation of a noisy sample $\mathbf{x}_{t}$ at timestep $t$. The LDM objective is then formulated as follows:
@@
Given the large success of LDMs, new mechanisms have appeared on how to control their image generation process further. In this paper, we use the ControlNet [55] as a backbone. The main idea behind ControlNet is to create a trainable copy of a pre-trained LDM, attach it to the LDM and in turn train it without modifying the original model (zero-convolutions). In this way, generated images retain the quality from the original LDM which remains unchanged. The copy is additionally conditioned on images $\mathbf{c}_{\text{f}}$ that are encoded into its latent space by a small learnt neural network and the feature maps of the encoder of the LDM copy are added to the feature maps of the original LDM so as to transfer the original conditioning signal. Including the text conditioning $\mathbf{c}_{t}$, the objective function for a conditional LDM is now equal to:
@@
We attach a ControlNet module to the Stable Diffusion model and train it on approx. $23K$ images, where the input conditioning are RGB HPM images and the output are the corresponding images of real people in the wild. During training, we use a generic text prompt that is dropped $50\%$ of the time. This model is later used as a translator for our custom low-resolution RGB HPM images that are padded and then upscaled from $32\times 64$ to $512\times 512$ during training. The output is upscaled from $200\times 200$ to $512\times 512$ during training.
Figure 3: Analysis of the StyleGAN generator: A. Contextual HPM Generation: Fixing the context (rows F) generates more uniform clothing than varying the context embeddings (rows R). This indicates that there is a link between the context and the environment that has been learned by our context-aware HPM generator.
B. Background Reconstruction: GT/VAE reconstructions are shown in rows 1/2; fixing the context vector in the StyleGAN generator gives matching results (row F) with GT/VAE, while varying the context embeddings leads to random predictions (row R).
@@ 3.4 Inpainting and super-resolution
The human body is highly complex due to its large range of poses and also appearances, as one body may be clothed in many different types of outfits. For this reason, there exists a large semantic gap between the body and its binary segmentation mask. This gap makes it especially difficult to inpaint human bodies into photographs, when only the ambiguous binary mask is available and the body is to be inpainted into the specified region. To this end, at inference time, we condition the inpainting LDM on a semantic representation of the body and clothing (HPM) via our pre-trained HPM translation module (ControlNet). In this way, we are able to transfer the background from the original image into the resulting image using the inpainting model while predicting the HPM-compatible humans at the same time. The impact of the HPM translator can be adjusted with a weight $\beta$ (controlNet strength) that is used for the multiplication of the ControlNet’s feature maps prior to their addition with the feature maps of the inpainting model. A low weight leads to a reduced guidance while producing higher quality imagery. We found that setting $\beta$ to $0.5$ leads to satisfactory results.
Table 1: Quantitative analysis: C denotes whether the context vector is fixed (F) or random (R). Fixing the context vector leads to predictions that are closer to GT, especially in background reconstruction. The results are calculated over 10K contexts. The analysis confirms that our method is context-aware (PSNR, MSE, LPIPS) and generates meaningful HPMs and backgrounds (FID, KID). The performance of the VAE module used to generate the blurry contexts (rows 2 and 5) is shown in row 7.
@@ C Evaluation PSNR $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
4 Experiments
Figure 4: Comparison against state of the art: we compare (A) StyleGAN-Human, (B) TextHuman, (C) ControlNet (OpenPose) & SD-2.1-Inpainting, with (D) Our full method. (A) is unconditional, while (B) is conditioned on a dense pose. Both do not respect the context - considering clothing semantics, and lighting. Parts of the body may be missing (see (B), col 2). While (C) is conditioned on a pose, the pose is not always respected. It changes the original scene (e.g. by insertion of new objects; see col 3) and it does not always generate single humans (see cols 1, 5). In contrast, Ours (D) does not change the scene, it generates single humans that are posed correctly, and overall contextual alignment, when clothing semantics and lighting are considered, is higher than in (A), (B), and (C). Note: (B), (C), and (D) have the same input pose.
4.1 Data
Our method utilizes the AHP dataset [61] comprising context images, images with humans, machine-annotated 2D pose information, and HPMs. For training, we extract context images by using segmentation masks to remove people from the images. Skeleton images with pose data are generated using OpenPose [7, 8]. During inference, only a context image and the pose information are required.
@@ 4.1 Data
The AHP dataset is compiled from large-scale instance segmentation and detection datasets, including COCO [31], VOC [11], SBD [17], and Objects365 [48]. It underwent filtering and post-processing to yield 56K images with parsing maps and binary human masks, which are integral to our method. However, the raw AHP dataset’s varied resolution and quality, along with inconsistent human pixel percentages, lead to sub-optimal results. Additionally, the provided parsing maps sometimes have disconnections and holes. To mitigate these issues, we further refined the AHP images. Images were selected based on human size, with a minimum dimension of $180$ pixels in height and 90 pixels in width, and standardized to a resolution of $200\times 200$. We also filtered out images with mismatched binary masks and parsing maps to prevent background bleeds into human images. To address occlusions in parsing maps, we added a segmentation category for accessories and discarded images with disconnected parsing map parts. Our final dataset consists of $\sim 23$K images, each paired with its corresponding parsing map and OpenPose keypoints, ensuring higher consistency and quality for training.
[…]
Output $\psi=0.5$ $\psi=0.6$ $\psi=0.7$ $\psi=0.8$ $\psi=0.9$ $\psi=1$
HPM 15.73 14.59 13.53 12.64 11.96 11.49
Context 112.45 86.33 67.55 54.77 46.29 41.08
[…]
Method $\beta=0$ $\beta=0.25$ $\beta=0.5$ $\beta=0.75$ $\beta=1$
Baseline - - 78.7 86.7 90.0
Ours 42.3 22.6 21.0 19.8 20.2
4.2 Comparison to state of the art
@@
Qualitative/quantitative analysis: Figure 4 contains a comparison with state-of-the-art full-body generation methods including (1) an unconditional GAN-based method (StyleGAN-Human [13]), (2) dense pose conditioned GAN-based method (TextHuman [21]) as well as (3) a 2D pose conditioned diffusion-based method (ControlNet [55]) using the same inpainting module [43] as in Ours. In (1), humans are generated at $1024\times 1024$, the background is subsequently removed using rembg [3] and humans are then pasted on top of the context image, at random. Next, in (2), humans are generated at $512\times 512$ from their randomized HPM predictions, and the same procedure is followed as in (1). In (3), rectangular masks are passed to the inpainting model as the 2D pose-conditioned ControlNet model does not provide information about the location of the body/clothing. Humans are inpainted at $512\times 512$. Similarly to (3), our method inpaints humans using a diffusion model, however, we use GAN-generated HPM masks for inpainting and additional guidance. Our method generates more natural-looking, environment-specific humans when the clothing semantics and the overall color harmonization are considered. Table 3 shows quantitative analysis against the 2D pose-conditioned ControlNet baseline. In Figure 5 (A), we gradually decrease the impact of our fine-tuned ControlNet model through $\beta$. When $\beta=0$, our system is equivalent to the vanilla Stable Diffusion inpainting model, using a binary mask (from the predicted HPM). It fails to inpaint humans into the photographs without additional guidance.
Figure 5: A. Super-Resolution: higher image quality may be achieved by lowering the ControlNet strength via $\beta$. However, when $\beta$ is too low, humans are no longer generated ($\beta=0$ shows the vanilla SD inpainting result without additional HPM guidance). B. Fixed pose: generated clothing changes depending on the context.
Figure 6: ESP outputs full-body humans wearing environment-inspired clothing that is semantically suitable for diverse contexts. We show a wide range of different scenes including indoors and outdoors, with varying image resolutions (from $200\times 200$ to $512\times 512$), target pose, and lighting conditions. The produced results depend on the quality of the context image and on the capabilities of the generic Stable Diffusion 2.1 inpainting model that is used in the final step of our system. Please zoom in for details.
Figure 7: Out-of-domain generalization: our results on top of synthetic backgrounds.
@@
Perceptual evaluation: To evaluate the appropriateness of our generated humans versus state-of-the-art, we use BLIP-2 [29] to caption a set of context images and group them by the environment type. We then select 20 prominent environment groups and randomly select one sample for each. We compare against the current state-of-the-art, i.e. a pose-conditioned LDM from [55]. 16 people attended the short study and the results show that our method was preferred 62% of the time. 3% of the time, the user was undecided.
@@ 4.3 Evaluation and ablation studies
Quantitative analysis: To determine the effectiveness of our HPM generator, we compare it to GT data with a fixed context C (denoted as F) and with a random context (R) in Table 1. The results demonstrate that our method generates meaningful parsing masks and background images when the original data distribution is considered (FID, KID). It is also shown that the background reconstruction errors (PSNR, MSE, LPIPS) are considerably lower when the context vector is fixed because our method generates context-aware predictions. The reconstructed backgrounds (row 2) are similar to the VAE predictions (row 7). The VAE module provides embeddings that allow for the HPM model to generate backgrounds that look similar to the real ones (row 3). Finally, Table 2 shows the ablation of the truncation trick $\psi$ parameter from  [25].
Context embeddings: Figure 3 (B) shows the blurry background reconstructions of our (1) VAE, and (2) contextual HPM StyleGAN generator. StyleGAN’s predictions are of lower quality than the VAE. Nevertheless, we find that these reconstructions are meaningful as they also lead to the generation of environment-inspired clothing semantics. In Figure 3 (A), the difference between the fixed and random context embeddings is shown in the generator. Fixing the context (rows F) generates more uniform clothing. Here, each sample is generated with a different random part of the contextual style vector, while the context embeddings are fixed. On the contrary, when the context embeddings are not fixed, our model generates more varied clothing styles (rows R). This indicates there is a link between the context and the environment.
Inpainting: Changing the impact of our ControlNet model $\beta$ may increase image quality (Figure 5 (A)). This is due to the low resolution of images that we use to train the model. When $\beta$ is lowered, the original Stable Diffusion model that was trained on high-quality images has more impact on the final result, and therefore images of higher quality are generated. However, if $\beta$ is too low, generated humans might no longer respect the input pose or they might not be generated, as discussed above in Sec. 4.2. For numerical analysis, see Table 3.
Pose conditioning: Previously, it had been shown that pose plays an important role in the synthesis of scene-compatible humans [6]. Here, we argue that the clothing semantics may also be inferred from the scene context alone. We fix the pose in Fig. 5(B) and Fig. 7 to show how our method generates plausible clothing. More examples with the original pose are shown in Fig. 1 and Fig. 6.
Limitations: Our method generates people in simple poses only (see Sec. 4), and, if the conditioning pose is not plausible, the method might fail to generate a valid HPM. Also, the quality of our results depends on the capabilities of the generic inpainting model and may be further enhanced using an upscaler (Fig. 8).
@@
Potential negative impact: Our system is biased as there is a lack of racial diversity in specific contexts, and the gender imbalances skewed toward male representations are a reflection of the training data’s limitations. Therefore, it is imperative to take conscious steps to mitigate these biases. This can include diversifying the training data which is out of the scope of this paper.
Figure 8: Upscaling: the results of our method (top row) may be further improved using an upscaler (bottom row) [1].
@@ 5 Conclusion
We have demonstrated how full-body human generation and subsequent inpainting can be made context-aware, i.e., our generated humans are in sync with the environment when their clothing semantics and overall color harmonization are considered. Our main contribution is conditioning the image generation process on scene context to generate environment-specific people in clothing that is suitable for inpainting into a given environment. Our results demonstrate that our method produces plausible contextualized people in scenes, outperforming state-of-the-art methods on the task of contextual full-body generation.
Acknowledgements: We thank Soubhik Sanyal for his assistance in writing, and Benjamin Pellkofer for IT support. This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska Curie grant agreement No 860768 (CLIPE project).